OpenStudy (anonymous):

Below are two expressions. Simplify each and choose the statement that is true about each. Expression #1 (4x)^2(x) Expression #2 (6x^2)^4 A. The exponents in Expression #1 are greater than the exponents of Expression #2. B. The exponents on Expression #2 are greater than the exponents of Expression #1. C. The exponents of Expression #1 are the same as the exponents of Expression #2. D. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

Please, anybody?? I need help! :(

OpenStudy (anonymous):

D - the values of the exponents depends on value of x so it cannot be determined

OpenStudy (anonymous):

Thanks for your help! The first equation confused me. I guess that's why I didn't understand...

OpenStudy (anonymous):

Do you think you could help me with this one as well?? Below are two expressions. Simplify each and then choose the statement that is true. Expression #1 (y^9)(2y^2)^3 Expression #2 (9y)(2y^3)^2 A. The exponents in Expression #1 are greater than the exponents of Expression #2. B. The exponents on Expression #2 are greater than the exponents of Expression #1. C. The exponents of Expression #1 are the same as the exponents of Expression #2. D. The relationship cannot be determined with the given information

OpenStudy (anonymous):

yw - if for example x = 5 then the exponent in first expression = 10 > no 2 and if x = 1 then its less

OpenStudy (anonymous):

I asked this question twice and the other person said this for the first one: B. The exponents on Expression #2 are greater than the exponents of Expression #1 since exponent in #2 is 2* 4 = 8 ( > 2 + 1 = 3)

OpenStudy (anonymous):

Which one is right? lol :)

OpenStudy (anonymous):

Renee99, you're confused my reply to the other problem !!!

OpenStudy (anonymous):

So it is B?

OpenStudy (anonymous):

(y^9)(2y^2)^3 = (y^9)(8y^6) = 8y^15 (9y)(2y^3)^2 = (9y(4y^6) = 36y^7 A

OpenStudy (anonymous):

Okay, thanks for writing that out. But i'm still confused on the first one. One of you said it was B and the other said it is D? Which one? lol :) I just want to be sure..

OpenStudy (anonymous):

Which one are your talking about?

OpenStudy (anonymous):

They first one... on the very top.

OpenStudy (anonymous):

Can you just copy and paste into here?

OpenStudy (anonymous):

Below are two expressions. Simplify each and choose the statement that is true about each. Expression #1 (4x)^2(x) Expression #2 (6x^2)^4 A. The exponents in Expression #1 are greater than the exponents of Expression #2. B. The exponents on Expression #2 are greater than the exponents of Expression #1. C. The exponents of Expression #1 are the same as the exponents of Expression #2. D. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

Now that's the one I have answered!

OpenStudy (anonymous):

Yeah you said it was B, but jimmyrep said it was D?

OpenStudy (anonymous):

x^2 * x = x^3 (x^2)^4 = x^8 -> B ( since 8 > 3)

OpenStudy (anonymous):

Do you understand the way I'm calculating?

OpenStudy (anonymous):

Okay, I'll trust you!! :D lol Thanks for your help!!

OpenStudy (anonymous):

Yes, I think I understand now!! :D

OpenStudy (anonymous):

Sometimes the screen jerking too much, causing typo a lot

OpenStudy (anonymous):

illl have another look at q 1 (x)^2(x) = x^2x (x^2)^4 = x^8 now if x = 1, 2nd is greater than ist if x = 5 1st is greater than 2nd so D

OpenStudy (anonymous):

The question asking about which EXPONENT, so I just focus on it

OpenStudy (anonymous):

i have assumed its x the the power (2x) - maybe its meant to be x^2 * x in which case B is correct - the way its written is somewhat ambiguous

OpenStudy (anonymous):

Okay, Its making much more sense to me now! :)

OpenStudy (anonymous):

Similarly: (y^9)(2y^2)^3 = 9 + ( 2*3) = 15 (9y)(2y^3)^2 = 1 + ( 3*2) = 7 => A ( 15 > 7)

OpenStudy (anonymous):

I do have one more question... If you guys wouldn't mind helping me with it??

OpenStudy (anonymous):

sorry - i gotta go

OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x^3y^4)^2

OpenStudy (anonymous):

Its all good, just thought I would ask!! Thanks for all of your help!! :D

OpenStudy (anonymous):

I'm trying to insert equation, but my PC and connection so staggering !

OpenStudy (anonymous):

Its all good, I just appreciate you helping me at all!! :)

OpenStudy (anonymous):

Go ahead, Renee!

OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x^3y^4)^2

OpenStudy (anonymous):

(3x^3y^4)^2 = [(3)(x^3)(y^4)]^2 = (3)^2*(x^3)^2*(y^4)^2 = 9*x^6*y^8 = 9 x^6 y^8 If everything inside the parentheses is multiplied together, all you need to do is raise every individual part to the outer exponent.

OpenStudy (anonymous):

(3x^3y^4)^2 Expand it out: (3x^3)^2 = 9x^6 (y^4)^2 = y^8

OpenStudy (anonymous):

That's what I got so far...

OpenStudy (anonymous):

Your's looks a lot simpler, lol :)

OpenStudy (anonymous):

Try to make it less complicated!

OpenStudy (anonymous):

Yeah, I tend to make it more complicated than it suppose to be... idk why, I just do it. lol

OpenStudy (anonymous):

You need to practice good habit. It sure benefit you in the long run for much tougher problems!

OpenStudy (anonymous):

I'm trying.... I'm getting better, slowly but surely. lol :D

OpenStudy (anonymous):

Can you help me with part 1 of that problem.

OpenStudy (anonymous):

You know a saying " All ways come to Rome"

OpenStudy (anonymous):

However of course we want the most shorcut way!

OpenStudy (anonymous):

Yeah, haha. Nice saying!! :)

OpenStudy (anonymous):

Explain Expand the expression, applying power rule!

OpenStudy (anonymous):

There're many way to solve one problem, but I'm different from others because I strive for the simplest and neat solution!

OpenStudy (anonymous):

The way my teacher explains it really breaks it down and it can sometimes make me confused. The way you did it made it seem so simple.

OpenStudy (anonymous):

For example, the goal of your question is exponent, so I just aim to solve for exponent!

OpenStudy (anonymous):

Yeah, Understandable!! :)

OpenStudy (anonymous):

one-by-one tutoring is the most efficient way!

OpenStudy (anonymous):

I'm home school, so It;s a little harder to get one on one time with an teacher, but I'm going to ask if there's a way to get extra help.

OpenStudy (anonymous):

home schooled*/ it's*

OpenStudy (anonymous):

I'm feel sorry for many askers who just want the answers for the purpose of score!

OpenStudy (anonymous):

How old are you? Are you taking algebra 1 now?

OpenStudy (anonymous):

I know, I see many kids who only like home schooling for the benefit of being able to get the answers through the internet.

OpenStudy (anonymous):

I'm 16 and in 10th grade. I know i'm taking it late but I had some issues last year and I ended up having to complete ALG1 this year.

OpenStudy (anonymous):

"Better late than never" :)

OpenStudy (anonymous):

Right!! :)

OpenStudy (anonymous):

Remember the story "Racing between the turtle and rabbit"?

OpenStudy (anonymous):

Yeah...

OpenStudy (anonymous):

I thought It was the tortoise and the hare, lol :P

OpenStudy (anonymous):

One of my cousin who 16, taking Advanced Calculus, just drop out,

OpenStudy (anonymous):

So you or she will win in the end :)

OpenStudy (anonymous):

Whoa, sorry to hear that but Damn, advanced calculus at 16. she must be smart

OpenStudy (anonymous):

Hahaha

OpenStudy (anonymous):

Most of her friends do it, All the stuff I'm showing you is 4th grade math in Asia!

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Are you serious??

OpenStudy (anonymous):

Now I feel like our country is so dumb.... lol

OpenStudy (anonymous):

jk :)

OpenStudy (anonymous):

Just point out to you that consistently patient is the key to success! Yep, Asian learn exponent operations very early!

OpenStudy (anonymous):

It's pointless if you win half of way, then neglect snobbish => Loser in no time!

OpenStudy (anonymous):

You got that right!! That's really shocking to hear. hahaha I always new Asians were smart, but I guess I never knew how smart. I do now, lol :) No wonder every kid I talk to nowadays seems so stupid and uneducated. What is happening to America. This generation is going to destroy the world! hahaha jk :)

OpenStudy (anonymous):

BTW I would just like to say again, Thanks for all of your help and that I really appreciate it!! :)

OpenStudy (anonymous):

The issue in this country is "liberty, freedom" !

OpenStudy (anonymous):

Exactly!!

OpenStudy (anonymous):

Heard that...

OpenStudy (anonymous):

Study for your future is an obligation, but here it considers boring, not fun! Interpret as your future will be not fun, too!

OpenStudy (anonymous):

You never go to regular school?

OpenStudy (anonymous):

Yeah I know, It's kind of sad... I take my schooling very seriously and I want to strive to be the best I can be!! But then I see kids who don't even care and it pisses me off! No, I went to public school most of my life, up until about three years ago. Once I reached middle school and high school, I absolutely hated it, but the home school program i'm on is really awesome! :)

OpenStudy (anonymous):

Good for you that you don't give up!

OpenStudy (anonymous):

Never!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Middle school is horrible and high school even worse!

OpenStudy (anonymous):

Keep in mind that I'm here for your academic success, so don't forget to take full advantage of it :)

OpenStudy (anonymous):

I know!! There's so much drama and craziness involved in the social life of teenagers its completely barbaric. I truly have no idea as to how it got so bad.

OpenStudy (anonymous):

I really appreciate the help everyone on here gives me! I plan to take full advantages in all opportunities that can help benefit me in school.

OpenStudy (anonymous):

As matter of fact, I'm surprised that you're NOT typical teenagers, stereotyped as rotten spoiled! Your parent mus be so proud, and delightful to have such a decent child as you!

OpenStudy (anonymous):

I'm actually helping several other high schoolers, however, they just dump all their homework on me, not like you trying to get to bottom of the concept.

OpenStudy (anonymous):

I'll let you come back to your homework now!

OpenStudy (anonymous):

Aww thanks!! It's funny, my parents tell me all the time that I'm "the golden child" or "the diamond in the rough" and that i'm not like most teenagers. Most of their friends that I meet even assume that I'm much older than I actually am and that I'm really mature or my age. I just look at it as rational thinking I guess....

OpenStudy (anonymous):

a good upbringing helps!! :)

OpenStudy (anonymous):

SEE, I can recognize it even just through online contact! Not really the environment, may be your family is super with religious

OpenStudy (anonymous):

Actually, that was the last of my homework!! :) Thanks again for all of your help. I really like how you broke it down for me instead of just giving me the answer like some people have in the past. I truly appreciate it!! :D

OpenStudy (anonymous):

That's just how I am, so slow to the degree that I'm always is the last to respond :P